Researchers in the corporate office of an airline wonder if there is a significant difference between the cost of a flight on Priceline.com vs. the airline's own website. A random sample of 16 flights were tracked on Priceline and the airline's website and the average difference between the vendors was $2.435 with a standard deviation of $7.6328. The 99% confidence paired-t interval for price (Priceline - Airline Site) was (-3.1879, 8.0579). Which of the following is the appropriate interpretation?

Question 4 options:

1)

The proportion of all flights that had a difference in price between the two vendors is 99%.

2)

We are 99% confident that the average difference in the prices of the flights sampled is between -3.1879 and 8.0579.

3)

We are certain the average difference in prices between the two vendors for all flights is between -3.1879 and 8.0579.

4)

We are 99% confident that the difference between the average price on Priceline and the average price on the airline's website is between -3.1879 and 8.0579.

5)

We are 99% confident that the average difference in price between the two vendors for all flights is between -3.1879 and 8.0579.

Question 5 (1 point)

The Department of Transportion (DOT) is attempting to determine the proportion of drivers who require all passengers in the car to wear their seatbelt before putting the vehicle in drive. A survey of 75 drivers is performed and 28 people say they will not drive until all passengers in the vehicle are buckled up. To report their finding they want to create a 90% confidence interval. What would be the margin error for this confidence interval?

Question 5 options:

1)

0.0559

2)

0.0916

3)

0.0051

4)

0.0715

5)

0.0106

Question 6 (1 point)

Saved

You work for the consumer insights department of a major big-box retailer and you are investigating the efficacy of a new e-mail marketing campaign. Through the use of e-mail analytics research, you have determined that in a random sample of 714 monitored subscribers, 181 of them opened the e-mail within 24 hours of receiving it. What is the 95% confidence interval for the true proportion of all e-mail subscribers that opened the e-mail within 24 hours of receiving it?

Question 6 options:

1)

( 0.23722 , 0.26978 )

2)

( 0.71459 , 0.77841 )

3)

( 0.22672 , 0.28028 )

4)

( 0.22159 , 0.28541 )

5)

( -0.22159 , 0.28541 )
Answer #1

4)

We are 99% confident that the average difference in price between the two vendors for all flights is between -3.1879 and 8.0579.

5)

sample success x = 28
sample size          n= 75
sample proportion p̂ =x/n= 0.3733
std error se= √(p*(1-p)/n) = 0.0559
for 90 % CI value of z= 1.645
margin of error E=z*std error   = 0.0916

6)

sample success x = 181
sample size          n= 714
sample proportion p̂ =x/n= 0.2535
std error se= √(p*(1-p)/n) = 0.0163
for 95 % CI value of z= 1.960
margin of error E=z*std error   = 0.0319
lower bound=p̂ -E                       = 0.22159
Upper bound=p̂ +E                     = 0.28541

option 4) ( 0.22159 , 0.28541 )